OpenStudy (anonymous):

Determine the indicated angle measure to the nearest degree drawing to follow

OpenStudy (anonymous):

 top line is 6 left line is 11.4 top right angle is a bottem angle is 31

OpenStudy (anonymous):

\[\left[\begin{matrix}\sin 31 & \sin a \\ 6.0 & 11.4\end{matrix}\right]\]

OpenStudy (anonymous):

is this correct ?